Consider the following reaction:

A+B+C→D

The rate law for this reaction is as follows:

Rate=k[A][C]2[B]1/2

Suppose the rate of the reaction at certain initial concentrations of A, B, and C is 1.19×10−2 mol⋅L−1⋅s−1 .

What is the rate of the reaction if the concentrations of A and C are doubled and the concentration of B is tripled?
Answer #1

Given that

Rate = k[A][C]2[B]1/2 = 1.19 *10-2 mol /L.s

Now, Both A and C are doubled where as B is tripled

SO, New rate = k [2A][3B]1/2[2C]2 = 8(3)1/2 k[A][C]2[B]1/2 = 8(3)1/2 ( 1.19 *10-2 mol /L.s) = 0.165mol/L.s or 1.65 *10-1 mol /L.s